Light travelling from water into crowd glass, strikes the crown glass at an angle of 41.7 degrees. What is the angle of refraction?
1
Expert's answer
2020-11-30T14:57:06-0500

We can find the angle of refraction from the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,"

here, "n_i=1.33" is the index of refraction of the incident medium (or the water), "n_r=1.52" is the index of refraction of the refractive medium (or the crown glass), "\theta_i=41.7^{\circ}" is the angle of incidence, "\theta_r" is the angle of refraction.

Then, from this equation we can calculate the angle of refraction:


"\theta_r=sin^{-1}(\dfrac{n_isin\theta_i}{n_r}),""\theta_r=sin^{-1}(\dfrac{1.33\cdot sin41.7^{\circ}}{1.52})=35.6^{\circ}"

Answer:

"\theta_r=35.6^{\circ}"